2 planes leave airport same time. 1 travels 40mph faster than the other. If they are 2000 miles apart after 5 hours find there average speeds?
Answer by mananth(16946)

Plane I x mph
Plane II x+ 40
..
They are going away from each other.

So add up their speed.

combined speed = x+ 1 x + 40
( 2 x + 40 )
Time = 5 hours
Distance = 2000 miles
Distance = speed * time
( 2 x 40 )* 5 = 2000
10 x + 200 = 2000
10 x = 2000 + -200
10 x = 1800
/ 10
x= 180 mph Plane I
Plane II 180 + 40 = 220 mph